# Creative Problem Solving in Large Language and Vision Models – What Would it Take?

### v2

@@ 1 Introduction
While recent state-of-the-art large language models (LLMs) and vision-language models (VLMs) have demonstrated competency in artistic endeavours Rombach et al. (2021); Copet et al. (2023), creative problem solving continues to be a shortcoming of these models (we use LLVM to denote the umbrella of both LLMs and VLMs). For instance, in Bubeck et al. (2023), the authors point out that “discontinuous tasks” that require a certain “Eureka” idea, i.e., creative problem solving, is currently a limitation of models like GPT-4. Similar observations have been made in follow up work showing that state-of-the-art LLMs inherently possess poor creative problem solving capabilities compared to humans Tian et al. (2023); Naeini et al. (2023). Given this obvious limitation, ongoing research in Machine Learning should seek to address the gap between LLVMs and creative problem solving, to further enhance the intelligent capabilities of these models. As defined in prior work, “Intelligence is the ability to work and adapt to the environment with insufficient knowledge and resources.Pennachin and Goertzel (2007), p.10. Demonstrated in hallmark examples of human ingenuity, like the makeshift $CO_{2}$ filter built onboard the Apollo-13 Cass (2005), or the makeshift medical devices used to offset equipment shortages during COVID-19 Turner et al. (2020), creative problem solving is especially important when dealing with resource-critical scenarios. Since humans may tend to “choke” under high pressure situations DeCaro et al. (2011) often limiting their CPS skills, autonomous agents equipped with LLVMs that have similar capabilities would be highly assistive and transformative to humans in high-stake environments. These include situations like rescue missions BBC (2012) or autonomous operation in human-inaccessible environments (e.g., space or underwater exploration) with limited resources Atkeson et al. (2018). However, the exceptional degree of creative problem solving necessary for such assistance remains beyond the scope of LLVMs today, limiting their intelligence (See Appx. B.1).
We believe that a discussion of Computational Creativity is essential to addressing this limitation. It is our position that Machine Learning and Computational Creativity should be strongly integrated in research to enable effective creative problem solving in LLVMs and push the frontiers of their ingenuity.
@@ 3 From Task Planning to Creative Problem Solving
Creative problem solving can be broadly described as the process through which agents discover novel ways of accomplishing a task that, prior to the discovery, was unsolvable. Computationally, creative problem solving can be achieved through planning, learning, or hybrid approaches Gizzi et al. (2022). Following a review of the different definitions of creative problem solving that have been proposed (Appx. A), we believe the following most closely connects to existing formalisms in ML.
@@ 3.1 Definition of Creative Problem Solving
As a simplified example, let us assume a robot that has a goal $G$ of transferring beans from a jar to a cooker: $G=${$in$(beans, cooker)}. Here, the initial state is defined as $C_{S}=${$in$(beans, jar), $hasContainability$(spoon)}. Let the actions be defined as $C_{A}=${$scoop$(beans, $X$, $loc_{s}$, $loc_{d}$)}, where, $X$ refers to an object that satisfies $hasContainability(\cdot)$ (e.g., spoon), to scoop beans from $loc_{s}$ to $loc_{d}$. If the robot has access to a spoon, the robot can use it to scoop the beans from the jar to the cooker. However, what if the robot did not have a spoon, but had a glass instead? By the definition of $C_{S}$, the agent is unaware that $hasContainability$(glass) is true, making the goal un-achievable. By our definition, creative problem solving is the process by which the agent uses some function $f(\cdot)$ to discover a new conceptual space: $f(C_{S})=C^{\prime}_{S}=C_{S}\mathop{\cup}${$hasContainability$ (glass)}. This would allow the agent to solve the previously unsolvable task by using the glass to scoop the beans instead.
Boden’s three forms of creativity denote three plausible functions for $f(C_{X})$. CPS arises when the agent uses what it knows, to discover something new and the newly discovered knowledge is applied to solve a previously impossible task. We revisit the notion of conceptual spaces in Section 3.
@@ 4 Augmenting LLVM embedding spaces for creative problem solving
In this section, we discuss how principles from CC can be extended to LLVMs for creative problem solving. We begin with Boden’s definition of “conceptual spaces” as “[conceptual space] is the generative system that underlies the domain and defines a certain range of possibilities: chess moves, or molecular structures, or jazz melodiesBoden (2005), p.18 and “… in short, any reasonably disciplined way of thinkingBoden (1998), p.214. By this definition, the embedding space of an LLVM describes its conceptual space or “its way of thinking”. Some evidence for this also comes from existing work that introduces an approach for enabling LLMs to interpret continuous embedding spaces via natural language. Given an embedding vector representing an interpolation of different concepts, the model is able to interpret a text prompt in the context of the supplied embedding Tennenholtz et al. (2023). The embedding thus determines the model’s way of thinking. Hence, a discussion of enabling creative problem solving in LLVMs should target their embedding space. To this end, we explore two questions: a) how can LLVM embedding spaces be augmented to achieve creative problem solving, and b) what information should they be augmented with? Aligning with our original position, we show that CC literature can offer insights into these questions.
@@ 4.1.3 Transformational Creativity
Transformational approaches involve transforming existing conceptual spaces to produce new ones. Transforming conceptual spaces can involve “altering existing rulesBoden (1998), p.216. One way of transforming a model’s embedding space involves fine-tuning or training Franceschelli and Musolesi (2023). However, additional insight into transformational creative problem solving comes from prior work in CC, that describes creative problems as those with a poorly defined structure where a solution is not immediately apparent Olteteanu (2014). And in such cases, “… re-representation being the process which transforms an ill-structured problem into a well-structured one with direct inference to a problem solutionOlteteanu (2014), p.1. The notion of “re-representing” or “redefining” the problem can be best captured in the input prompts provided to an LLVM. This most closely connects to prompt engineering and in-context learning (ICL).
Prompt engineering augments LLVMs with task specific hints, called prompts, to adapt the LLVM to new tasks Gu et al. (2023). Relatedly, in-context learning is a prompting method that provides the LLVM with instructions for solving a new task without requiring additional training. Prior work has shown that in-context learning and gradient-based optimization are equivalent Von Oswald et al. (2023), thus connecting ICL to training or fine-tuning.
Extension of transformational creativity to LLVMs: Task re-representations for creative problem solving, through prompting or ICL, has not been well explored within ML. Prompt engineering and ICL is a challenging task, since model performance depends strongly on the chosen prompts Rubin et al. (2021), further compounded by the fact that creative problems are inherently poorly defined Olteteanu (2014). However, useful insights can be derived from CC literature. For instance, regarding problems that require creatively re-purposing objects, the Object-replacement-object-composition (OROC) framework Olteţeanu and Falomir (2016) illustrates re-representations of tasks, that can be translated into prompts. The paper defines three different types of creative tasks involving objects, and their task re-representations as (from Olteţeanu and Falomir (2016), p.16):
@@
    Decompose objects. “If I do not have object $X$ with affordance $Af_{X}$ , which objects $Y_{1};Y_{2};...;Y_{n}$ which are components of object $Y$ could I use to obtain an object $Y^{\prime}_{i}$ with an equivalent or similar affordance, $Af_{X}\approx Af_{Y^{\prime}i}$?
For task re-representation, affordances can refer to object properties that are relevant to the task, e.g., in some cases the shape may be relevant and in other cases, the material Olteţeanu and Falomir (2016). Within LLVMs, the affordances $Af_{X}$ or $Af_{Y}$ can be defined via natural language, or other modalities such as images. In the following section, we present preliminary experiments on using LLVMs for object replacement, with prompts that are inspired by the above task re-representations. However, an in-depth application of these re-representations as defined in CC to in-context learning in LLVMs remains unexplored.
@@ 4.2 What information should LLVM embeddings be augemented with?
In the previous section, we discussed three methods for augmenting LLVM embedding spaces. In this section, we explore the question: “What information should be targeted by the three methods when augmenting the embedding space for creative problem solving?”. In the previous section, we discussed this in the context of OROC. According to the OROC framework Olteţeanu and Falomir (2016), information about object affordances could enable models to re-represent the task, such that the solution becomes evident. We propose a small experiment to validate whether the principles of transformational creativity from OROC are useful to LLVMs. We note that creativity can occur in various contexts, e.g., creatively solving a math problem or creatively playing a chess move, each of which would require different information. However, to facilitate the discussion in this paper, we focus our scope on tasks that require innovatively replacing missing objects (OROC Task #1).
@@ 4.2.1 Experiment Setup
We create a simple experiment setup that tests the “object replacement” principle from OROC, where we create test sets composed of images of objects for replacing one of five core objects: “Scoop”, “Hammer”, “Spatula”, “Toothpick”, and “Pliers”. We create two groups of tests: a) a nominal group where the actual object itself is available in each test set and requires no replacement (which serves as a form of baseline), and b) an object replacement group, where the nominal tool is missing and a creative replacement object should be chosen.
For each group, we create test sets with 4 objects each, chosen from a set of RGB images of 16 objects (Appendix Figure 3). We create 10 such test sets per core object (total 50 samples per model). Each test set only includes one ground truth object, along with three other random objects that will not suit as an appropriate replacement. In the nominal group, the ground truth is the actual object itself. In the object replacement group, the replacements are chosen based on self-assessment of the authors as (core object $\xrightarrow{}$ replacement): “Scoop” $\xrightarrow{}$ “Bowl”; “Hammer” $\xrightarrow{}$ “Saucepan”; “Spatula” $\xrightarrow{}$ “Knife”; “Toothpick” $\xrightarrow{}$ “Safety pin”; “Pliers” $\xrightarrow{}$ “Scissors”. For each test case, we pass the images in the test set along with a prompt. We record whether the ground truth object image was chosen by the model for the prompt (i.e., assigned highest output probability) [^2] CLIP generates probabilities that given images correspond to a text. ViLT and LLaVA respond with a text, and we evaluate if the model responded “yes” with a high probability for the ground truth..
@@
    Baseline (regular) prompt: Same prompt as used in the nominal cases to obtain a baseline.
@@ 4.2.3 Summary
The models used in our experiments have all been trained jointly in visual and text domains. Multi-modal prompting capabilities may be useful for achieving creative problem solving. It can be quite challenging to describe affordances in words (example of “hammers” in our tests) and they may be better described through other means, e.g., images or depth maps or spectral data for material properties Erickson et al. (2020). This would require application of multi-modal LLVMs that can process a variety of data types Girdhar et al. (2023); Han et al. (2023). Computational creativity can offer insights into meaningful representations of these different modalities that would help achieve creative problem solving, e.g., whether object material or shape matters more for one task vs. another Olteţeanu and Falomir (2016).
It is also worth noting that the creative problem solving examples in our experiments are human-centric. For instance, robots may not have similar capabilities as humans to manipulate bowls for scooping. In such cases, LLVMs need to account for the affordances as described with respect to the agent, in order to derive creative solutions. However, that adds another level of complexity, yet to be explored, since these models are typically trained on human-centric data.
@@ 5 Evaluation of Creativity
An important discussion in the context of creative problem solving is, how can creative problem solving be evaluated?. Prior work has proposed that creativity necessitates both novelty and value Boden (1998); Runco and Jaeger (2012), where the former guarantees that the generated outputs of a creative process are original, and the latter ensures that the generated outputs are useful. In the context of CPS, novelty refers to the discovery of new concepts (as defined in section 3.1), whereas value insists that the newly discovered concepts successfully solve the task. Hence, benchmarks for CPS should specifically evaluate how the task was solved (novelty and value) rather than the typical ML evaluation of whether the task was successful or not (value only). Some existing approaches that make this distinction describe problem settings that can be used to measure CPS skills of LLMs through the implicit integration of novelty and value measurements Tian et al. (2023); Naeini et al. (2023); Bisk et al. (2020); Talmor et al. (2022). In Tian et al. (2023), the authors create a dataset of 1600 real-world problems that necessarily involve creative reasoning abilities. Their proposed benchmark involves identifying novel approaches that can accomplish the given task (value). Similarly, in Naeini et al. (2023), the authors introduce the Only-Connect-Wall (OCW) dataset to measure CPS capabilities of LLMs. The authors in Bisk et al. (2020) explore physical commonsense reasoning that is more generally applicable, beyond object-based creative problems. The authors introduce Physical Interaction: Question Answering, or PIQA consisting of 16,000 QA pairs where each question is paired with two possible common-sense solutions with a ground truth. In Talmor et al. (2022), the authors introduce CommonSenseQA 2.0 (CSQA2) dataset consisting of both object-based and non-object based creative problems. The dataset consists of 14,343 questions distributed across 1,868 distinct topics. Currently, to the best of our knowledge, there are no standard benchmarks available to measure CPS skills of VLMs, although our preliminary experiments show one way to measure this using the task of object substitution.
@@ 6 Conclusion and Future Work
In this paper, we argued that an effective approach for enabling creative problem solving – currently a key limitation of LLVMs – should derive from Computational Creativity literature. To emphasize this at each juncture, we discussed the specific principles from CC that can be extended to achieve creative problem solving in LLVMs, describing the potential for further research with these insights. It is rare to see special tracks or workshops targeted at Computational Creativity within more prestigious ML conferences. These programs typically focus on creative artifact generation and art (such as the NeurIPS Workshop on Machine Learning for Creativity and Design NeurIPS (2022) or the recent tutorial at EMNLP on Creative Natural Language Generation Chakrabarty et al. (2023)), but do not discuss CPS, thus failing to bridge the gap between CC and ML. We hope to see a deeper integration of the CC communities at such strong ML venues. We hope to encourage the reader to view creative problem solving and ML holistically, through the lens of Computational Creativity.

### v3

@@ 1 Introduction
While recent state-of-the-art large language models (LLMs) and vision-language models (VLMs) have demonstrated competency in artistic endeavours Rombach et al. (2021); Copet et al. (2023), creative problem solving continues to be a shortcoming of these models (we use LLVM to denote the umbrella of both LLMs and VLMs). For instance, in Bubeck et al. (2023), the authors point out that “discontinuous tasks” that require a certain “Eureka” idea, i.e., creative problem solving, is currently a limitation of models like GPT-4. Similar observations have been made in follow up work showing that state-of-the-art LLMs inherently possess poor creative problem solving capabilities compared to humans Tian et al. (2023); Naeini et al. (2023).
Given this obvious limitation, ongoing research in Machine Learning should seek to address the gap between LLVMs and creative problem solving, to further enhance the intelligent capabilities of these models. We believe that a discussion of Computational Creativity is essential to addressing this limitation. It is our position that Machine Learning and Computational Creativity should be strongly integrated in research to enable effective creative problem solving in LLVMs and push the frontiers of their ingenuity.
Creative problem solving can be a resourceful skill for artificial agents. As defined in prior work, “Intelligence is the ability to work and adapt to the environment with insufficient knowledge and resources.Pennachin and Goertzel (2007), p.10. Demonstrated in hallmark examples of human ingenuity, like the makeshift $CO_{2}$ filter built onboard the Apollo-13 Cass (2005), or the makeshift medical devices used to offset equipment shortages during COVID-19 Turner et al. (2020), creative problem solving is especially important when dealing with resource-critical scenarios. Since humans tend to “choke” under high pressure situations DeCaro et al. (2011) often limiting their problem solving skills, autonomous agents equipped with LLVMs that have similar capabilities would be highly assistive and transformative to humans in high-stake environments. These include situations like rescue missions BBC (2012) or autonomous operation in human-inaccessible environments (e.g., space or underwater exploration) with limited resources Atkeson et al. (2018). However, the exceptional degree of creative problem solving necessary for such assistance remains beyond the scope of LLVMs today, limiting their intelligence (See Appx. B.1).
@@ 3 From Task Planning to Creative Problem Solving
Creative problem solving can be broadly described as the process through which agents discover novel ways of accomplishing a task that was unsolvable prior to the discovery. Computationally, creative problem solving can be achieved through planning, learning, or hybrid approaches Gizzi et al. (2022). Following a review of the different definitions of creative problem solving that have been proposed (Appx. A), we believe the following most closely connects to existing formalisms in ML.
@@ 3.1 Definition of Creative Problem Solving
As a simplified example, let us assume a robot that has a goal $G$ of transferring beans from a jar to a cooker: $G=${$in$(beans, cooker)}. Here, the initial state is defined as $C_{S}=${$in$(beans, jar), $hasContainability$(spoon)}. Let the actions be defined as $C_{A}=${$scoop$(beans, $X$, $loc_{s}$, $loc_{d}$)}. Here, $X$ refers to an object that satisfies $hasContainability(\cdot)$, for example a spoon, to scoop beans from $loc_{s}$ to $loc_{d}$. If the robot has access to a spoon, the robot can use it to scoop the beans from the jar to the cooker to meet the goal. However, what if the robot did not have a spoon, but had a glass instead? By the definition of $C_{S}$, the agent is unaware that $hasContainability$(glass) is true, making the goal un-achievable. By our definition, creative problem solving is the process by which the agent uses some function $f(\cdot)$ to discover a new conceptual space: $f(C_{S})=C^{\prime}_{S}=C_{S}\mathop{\cup}${$hasContainability$ (glass)}. This would allow the agent to solve the previously unsolvable task by using the glass to scoop the beans instead.
In essence, CPS arises when the agent uses what it knows, to discover something new and the newly discovered knowledge is applied to solve a previously impossible task. Boden’s three forms of creativity that we focus on in this paper, denote three plausible functions for $f(C_{X})$. We revisit the notion of conceptual spaces in Section 3.
@@ 4 Augmenting LLVM embedding spaces for creative problem solving
In this section, we discuss how principles from CC can be extended to LLVMs for creative problem solving. We begin with Boden’s definition of “conceptual spaces” as “[conceptual space] is the generative system that underlies the domain and defines a certain range of possibilities: chess moves, or molecular structures, or jazz melodiesBoden (2005), p.18 and “… in short, any reasonably disciplined way of thinkingBoden (1998), p.214. By this definition, the embedding space of an LLVM describes its conceptual space or “its way of thinking. Some evidence for this also comes from existing work that introduces an approach for enabling LLMs to interpret continuous embedding spaces via natural language. Given an embedding vector representing an interpolation of different concepts, the model is able to interpret a text prompt in the context of the supplied embedding Tennenholtz et al. (2023). The embedding thus determines the model’s way of thinking. Hence, a discussion of enabling creative problem solving in LLVMs should target their embedding space. To this end, we explore two questions: a) how can LLVM embedding spaces be augmented to achieve creative problem solving, and b) what information should they be augmented with? Aligning with our original position, we show that CC literature can offer insights into these questions.
@@ 4.1.3 Transformational Creativity
Transformational approaches involve transforming existing conceptual spaces to produce new ones. Transforming conceptual spaces can involve “altering existing rulesBoden (1998), p.216. One way of transforming the embedding space of a model involves fine-tuning or training Franceschelli and Musolesi (2023). However, additional insight into transformational creative problem solving comes from prior work in CC, which describes creative problems as those with a poorly defined structure where a solution is not immediately apparent Olteteanu (2014). And in such cases, “… re-representation being the process which transforms an ill-structured problem into a well-structured one with direct inference to a problem solutionOlteteanu (2014), p.1. The notion of “re-representing” or “redefining” the problem can be best captured in the input prompts provided to an LLVM. This most closely connects to prompt engineering and in-context learning (ICL).
Prompt engineering augments LLVMs with task-specific hints, called prompts, to adapt the LLVM to new tasks Gu et al. (2023). Relatedly, in-context learning is a prompting method that provides the LLVM with instructions to solve a new task without requiring additional training. Previous work has shown that in-context learning and gradient-based optimization are equivalent Von Oswald et al. (2023), thus connecting ICL to training or fine-tuning.
Extension of transformational creativity to LLVMs: Task re-representations for creative problem solving, through prompting or ICL, have not been well explored within ML. Prompt engineering and ICL is a challenging task, since model performance is strongly dependent on the chosen prompts Rubin et al. (2021), further compounded by the fact that creative problems are inherently poorly defined Olteteanu (2014). However, useful insights can be derived from CC literature. For instance, regarding problems that require creatively re-purposing objects, the Object-replacement-object-composition (OROC) framework Olteţeanu and Falomir (2016) illustrates re-representations of tasks, which can be translated into prompts. The paper defines three different types of creative tasks involving objects, and their task re-representations as (from Olteţeanu and Falomir (2016), p.16):
@@
    Decompose objects. “If I do not have an object $X$ with affordance $Af_{X}$ , which objects $Y_{1};Y_{2};...;Y_{n}$ which are components of object $Y$ could I use to obtain an object $Y^{\prime}_{i}$ with an equivalent or similar affordance, $Af_{X}\approx Af_{Y^{\prime}i}$?
For task re-representation, affordances can refer to object properties that are relevant to the task, e.g., in some cases the shape may be relevant and in other cases, the material Olteţeanu and Falomir (2016). Within LLVMs, the affordances $Af_{X}$ or $Af_{Y}$ can be defined via natural language or other modalities such as images. In the following section, we present preliminary experiments on using LLVMs for object replacement, with prompts that are inspired by the above task re-representations. However, an in-depth application of these re-representations as defined in CC to in-context learning in LLVMs remains unexplored.
@@ 4.2 What information should LLVM embeddings be augemented with?
In the previous section, we discussed three methods for augmenting LLVM embedding spaces. In this section, we explore the question: “What information should be targeted by the three methods when augmenting the embedding space for creative problem solving?”. In the previous section, we discussed this in the context of OROC. According to the OROC framework Olteţeanu and Falomir (2016), information about object affordances could enable models to re-represent the task, such that the solution becomes evident. We propose a small experiment to validate whether the principles of transformational creativity of OROC are useful to LLVMs. We note that creativity can occur in various contexts, e.g., creatively solving a math problem or creatively playing a chess move, each of which would require different information. However, to facilitate the discussion in this paper, we focus our scope on tasks that require innovatively replacing missing objects (OROC Task #1).
@@ 4.2.1 Experiment Setup
We create a simple experiment setup that tests the “object replacement” principle from OROC, where we create test sets composed of images of objects for replacing one of five core objects: “Scoop”, “Hammer”, “Spatula”, “Toothpick”, and “Pliers”. We create two groups of tests: a) a nominal group where the actual object itself is available in each test set and requires no replacement (which serves as a form of baseline) and b) an object replacement group, where the nominal tool is missing, and a creative replacement object should be chosen.
For each group, we create test sets with 4 objects each, chosen from a set of RGB images of 16 objects (Appendix Figure 3). We create 10 test sets per core object (total 50 samples per model). Each test set only includes one ground truth object, along with three other random objects that will not suit as an appropriate replacement. In the nominal group, the ground truth is the actual object itself. In the object replacement group, the replacements are chosen based on self-assessment of the authors as (core object $\xrightarrow{}$ replacement): “Scoop” $\xrightarrow{}$ “Bowl”; “Hammer” $\xrightarrow{}$ “Saucepan”; “Spatula” $\xrightarrow{}$ “Knife”; “Toothpick” $\xrightarrow{}$ “Safety pin”; “Pliers” $\xrightarrow{}$ “Scissors”. For each test case, we pass the images in the test set along with a prompt. We record whether the ground-truth object image was chosen by the model for the prompt (i.e., assigned highest output probability) [^2] CLIP generates probabilities that given images correspond to a text. ViLT and LLaVA respond with a text, and we assess if the model responded “yes” with a high probability for the ground truth..
@@
    Baseline (regular) prompt: the same prompt as used in the nominal cases to obtain a baseline.
@@ 4.2.3 Summary
The models used in our experiments have all been trained jointly in visual and text domains. Multi-modal prompting capabilities may be useful for achieving creative problem solving. It can be quite challenging to describe affordances in words (example of “hammers” in our tests) and they may be better described through other means, e.g., images or depth maps or spectral data for material properties Erickson et al. (2020). This would require the application of multi-modal LLVMs that can process a variety of data types Girdhar et al. (2023); Han et al. (2023). Computational creativity can offer insight into meaningful representations of these different modalities that would help achieve creative problem solving, e.g., whether object material or shape matters more for one task vs. another Olteţeanu and Falomir (2016).
It is also worth noting that the creative problem solving examples in our experiments are human-centric. For instance, robots may not have similar capabilities as humans to manipulate bowls for scooping. In such cases, LLVMs need to account for the affordances as described with respect to the agent, to derive creative solutions. However, that adds another level of complexity, yet to be explored, since these models are typically trained on human-centric data.
@@ 5 Evaluation of Creativity
An important discussion in the context of creative problem solving is how can creative problem solving be evaluated?. Prior work has proposed that creativity necessitates both novelty and value Boden (1998); Runco and Jaeger (2012), where the former guarantees that the generated outputs of a creative process are original, and the latter ensures that the generated outputs are useful. In the context of CPS, novelty refers to the discovery of new concepts (as defined in section 3.1), while value insists that the newly discovered concepts successfully solve the task. Hence, CPS benchmarks should specifically evaluate how the task was solved (novelty and value) rather than the typical ML evaluation of whether the task was successful or not (value only). Some existing approaches that make this distinction describe problem settings that can be used to measure CPS skills of LLMs through the implicit integration of novelty and value measurements Tian et al. (2023); Naeini et al. (2023); Bisk et al. (2020); Talmor et al. (2022). In Tian et al. (2023), the authors create a dataset of 1600 real-world problems that necessarily involve creative reasoning abilities. Their proposed benchmark involves identifying novel approaches that can accomplish the given task (value). Similarly, in Naeini et al. (2023), the authors introduce the Only-Connect-Wall (OCW) dataset to measure CPS capabilities of LLMs. The authors in Bisk et al. (2020) explore physical commonsense reasoning that is more generally applicable, beyond object-based creative problems. The authors introduce Physical Interaction: Question Answering, or PIQA consisting of 16,000 QA pairs where each question is paired with two possible common-sense solutions with a ground truth. In Talmor et al. (2022), the authors introduce CommonSenseQA 2.0 (CSQA2) dataset consisting of both object-based and non-object based creative problems. The dataset consists of 14,343 questions distributed across 1,868 distinct topics. Currently, to the best of our knowledge, there are no standard benchmarks available to measure CPS skills of VLMs, although our preliminary experiments show one way to measure this using the task of object substitution.
@@ 6 Conclusion and Future Work
In this paper, we argued that an effective approach for enabling creative problem solving – currently a key limitation of LLVMs – should derive from Computational Creativity literature. To emphasize this at each juncture, we discussed the specific principles from CC that can be extended to achieve creative problem solving in LLVMs, describing the potential for further research with these insights.
It is rare to see special tracks or workshops targeted at Computational Creativity within more prestigious ML conferences. These programs typically focus on creative artifact generation and art (such as the NeurIPS Workshop on Machine Learning for Creativity and Design NeurIPS (2022) or the recent tutorial at EMNLP on Creative Natural Language Generation Chakrabarty et al. (2023)), but do not discuss CPS, thus failing to bridge the gap between CC and ML. We hope to see a deeper integration of the CC communities at such strong ML venues. We hope to encourage the reader to view creative problem solving and ML holistically, through the lens of Computational Creativity.